solve for the ratio u:v if uČ+3uv-10vČ=0
Found 2 solutions by Theo, Shin123:
Answer by Theo(13342)

factor the equation to get:

(u - 2v) * (u + 5v) = 0

solve for u to get:

u = 2v or u = -5v

divide both sides of these equations by v to get:

u/v = 2 or u/v = -5

i've never done one like this before, but i think that might be your answer.

i did some checking on the original equation and found that the equation is true when you assume these ratios.

from u/v = 2, i re-derived u = 2v.
i then let v = 2 and u = 4 and the original equation became true with those values for u and v.

from u/v= -5, i let v = 10 and u = -50 and the original equation became true with those values for u and v.

consequently, i believe your solution is that u/v = 2 or u/v = -5.









Answer by Shin123(626)

Factor to get
(u-2v)(u+5v)=0
u=2v or u=-5v
u/v=2 or u/v=-5 I checked with u=10 and v=5 for the first, and I checked with u=-10 and v=-2 for the second.